# An object-centric core metamodel for IoT-enhanced event logs

### v1

@@ 6.3 Data Transformation Layer
The Data Transformation Layer focuses on transforming the CORE data structures into the OCEL 2.0 standard. This transformation step is critical, as it operationalizes the conceptual design (5) by aligning the internal semantics of the CORE objects, events, and relationships with the structural and syntactic constraints of OCEL 2.0. At its core, this layer transforms the CORE metamodel classes (Figure 6) into OCEL classes (Figure 2). This transformation is performed in three phases: The first phase transforms physical items, digital artifacts, and data sources from the CORE metamodel into OCEL-compatible objects. As illustrated in Figure 12, the CORE implementation provides multiple utility classes. As described in Section 5.2.2, Core Objects are grouped into three categories (i.e., DataSources, BusinessObject, GeneralObject). This subdivision can be refined using the object_class attribute of the object class. Moreover, the enum contains distinct possible object classes that are semantically connected to categorize. For example, a sensor is likely to be a data source. In addition to the object_class attributes, the object_type and object_id attributes are assigned directly to the corresponding OCEL counterparts. Additionally, all use case-specific attributes are assigned in the columns ‘ocel:attr:[key]’, where the key represents the corresponding attribute label. Finally, each object_id must be unique; otherwise, the object will not be transferred.
@@
The second phase handles the event transformation. For that, the implementation provides an abstract Event class with three implementations (see Section 5.2.1, Figure 13). The event class requires four attributes, while the ProcessEvent implementation requires one additional activity attribute. The event_id, timestamp, event_class and attributes are mapped accordingly to the object implementation. The event_type is dependent on the event_class. First, if the event_class is process_event the event_type is the activity label. Next, if the event_class is iot_event the event_type is some meaningful label. Finally, if the event_class is observation, the event_type is hardcoded to observed.
@@
The last phase establishes and maintains three relationships between the entities. As previously outlined, utility classes are also present in this phase (Figure 14). They define the required attributes for the different relationships. Generally, the (event/object)_id is utilized for the mapping between the different entities. Each relationship is associated with a relationship_class, which is also serialized in the relationship table. As there is a direct equivalent in the OCEL 2.0 standard for the e2o and o2o relationships, these can be transferred directly. The only exception is the mandatory activity that must be used for each relationship (in OCEL). As this is not required for every event_class, the event_type is used in those cases. On the other hand, the event-to-event relationships require more parsing because there is not a corresponding definition in the OCEL standard. Given that, a linking object without semantically relevance is utilized. Therefore, two relationships are established. One between the first event and the linking object, and another between the linking object and the second event. Hence, the data can be queried as there are directly connected.
@@ 6.4 OCEL Representation Layer
The implementation should be independent of a specific OCEL implementation (e.g., pm4py.OCEL [(Berti et al. 2023)], ocpa [(Adams et al. 2022)]) and follow the structure provided in Figure 2. This abstraction not only enhances the maintainability and flexibility of the implementation, but also ensures that the output remains reusable and interoperable.
@@ 7.1.1 Use case description
In this use case, we describe how we transcribe a log originally stored in the NICE format to our common CORE metamodel . This synthetic event log, first presented in (Bertrand et al. 2023), represents the behavior of two inhabitants living in a smart home for four weeks. To generate it, we used the smart spaces simulator presented in (Veneruso et al. 2023), which we adapted to generate NICE logs. More specifically, the two users were programmed to execute daily life activities such as Cook and eat, Do the dishes, Drink, Eat cold, Eat warm, Exercise, Go work, Sleep, Rest, Shop, Use Computer, Watch tv and Wc. Moreover, the simulator was programmed so that during one week, the users show a different behavior, i.e., they stay at home the whole day instead of going to work in the morning and coming back home in the evening during weekdays. This change in behavior is due to extremely high outside temperatures during that week, which are tracked by temperature sensors and logged as IoT events.
@@
  • Flexibility of the process: human routines are characterized by their context dependency, their lack of structure, the repetition of activities and the execution of multiple activities in parallel (Leotta et al. 2015);
@@ 7.2.1 Use Case Description
In manufacturing processes, multiple (IoT) devices operate collaboratively to produce a part. These processes encompass various steps, including the handling of parts, which may involve multiple machines, transportation between different machines, and the measuring of parts. In the following, two use cases are presented that demonstrate the integration of IoT devices into manufacturing processes. In each case, the processes are managed by a process execution engine and are logged using the XES DataStream extension, ensuring real-time tracking and analysis of production events. This approach facilitates enhanced process monitoring, optimization, and traceability, contributing to improved efficiency and product quality. (Mangler et al. 2023):
  • In the IoT lab at Trier University [^4] More information can be found at https://iot.uni-trier.de., a Fischertechnik learning factory is used to carry out research related to BPM (Malburg et al. 2020). The use case includes machines which are equipped with several types of sensors that collect production-related data, such as the state of connected light barriers or switches. Complementing this Internet of Things (IoT) data, information pertaining to process executions is systematically recorded and maintained as event logs, facilitating comprehensive monitoring and analysis of manufacturing processes (Grüger et al. 2023).
@@ 7.2.2 Translation to core log format
The logs for the use cases described before are logged using the XES DataStream extension (Mangler et al. 2023). In the following, it is described how they can be translated into the core model format:
@@ 7.3.1 Use case description
In this use case, the focus is on the IoT-supported online monitoring and conformance checking of the blood donation process, augmented with hand hygiene guidelines from the World Health Organization (WHO) that prescribe additional process steps of hand disinfection (Franceschetti et al. 2023a). The goal is to be able to provide a timely feedback to the healthcare workers regarding the conformance with the hand hygiene requirements, reducing the risk of contamination that could threaten the patients’ health. As the process is fully manual and executed without the support of a process engine, the challenge here is to be able to generate in real-time IoT events from the process execution and abstract them into the respective process events. However, ambiguities deriving from multiple process events such as Apply tourniquet and Insert needle being represented by the same IoT events pose a challenge to the event abstraction. This results in multiple process events being generated from a given set of IoT events, with attached a quantification of the respective ambiguity.
@@ 8 Discussion
The set of requirements presented in Sect. 4, being derived from the harmonization of requirements from different models addressing diverse use cases, provides a solid foundation for the development of a suitable representation of IoT-enhanced process logs. The proposed CORE metamodel successfully fulfills these requirements, as reported in Sect. 7. Specifically, the metamodel effectively supports different use cases in various IoT domains. Moreover, the proof-of-concept Python implementation and its application to concrete use cases underscored the suitability of the metamodel for multiple near real-world applications.
The CORE metamodel offers the following major advantages for researchers:
  • Unified Representation: The model provides a standardized approach to integrating IoT data, process data, and contextual information across different levels of granularity. This offers common reference points for future research, facilitating cross-study comparisons, meta-analyses, and the development of new process mining algorithms.
  • Flexibility and Extensibility: By supporting semantic annotations, metadata representation, and diverse data sources, the model accommodates a wide range of use cases, as demonstrated in the evaluation in Sect. 7. This extensible architecture paves the way for future extensions and customized adaptations, for example capturing domain-specific knowledge, thereby driving research in semantic process mining and context-aware analytics.
  • Interoperability: The choice of implementing the CORE metamodel within the OCEL 2.0 format facilitates the uptake by researchers and the integration with existing use cases, fostering data exchange and reuse.
@@
Moreover, the CORE metamodel yields the following implications for practitioners:
  • Data Exchange: By adopting the CORE metamodel , organizations can translate and integrate data from various sensors, information systems, and legacy event log formats. This facilitates data exchange between different systems, increasing the opportunities to engage with process mining analyses.
  • Facilitated Process and IoT Integration: The metamodel design allows for the simultaneous representation of process events and IoT observations. This duality enables practitioners to bridge the gap between the physical world and digital process representations, thus enhancing the ability to monitor and analyze business processes.
@@
Overall, the CORE metamodel addresses several challenges, providing benefits for researchers and practitioners. Still, we recognize some limitations of the metamodel. Specifically, the reconciliation of diverse requirements from different use cases and research groups resulted in a number of trade-offs in the expressive power of the metamodel. This results in the lack of explicit concepts in the CORE metamodel able to represent use case-specific concepts, such as the Formal Model concept from CAIRO, which represents a formal process specification and is relevant in the smart healthcare use case. Nevertheless, the CORE metamodel remains open for specialization and extension to ensure that also use-case specific requirements can be met. This can, for example, be achieved by leveraging event attributes, or by specializing the general object concept and its associations. We emphasize that this reflects deliberate design choices made during the metamodel development to prioritize generalizability and interoperability–which we consider major goals of the metamodel.
@@ 9 Conclusion and Future Work
In conclusion, the contribution of this paper is manifold: 1) we present a set of requirements, important for the development of a suitable representation format for IoT-enhanced event logs, 2) we propose a new model referred to as CORE metamodel for IoT-enhanced event log, rooted in the object-centric paradigm to foster data exchange in the community, 3) we present our prototypical Python implementation of this new model, compatible with the PM4PY package and the OCEL 2.0 specification, and 4) we use this implementation for evaluation in several near real-world IoT use cases. To ensure the quality of the model and rigor in its design, the model has been created following the design science research methodology, building on existing IoT ontologies and IoT-enhanced log models, and based on a set of requirements gathered from experts in IoT-enhanced PM. Furthermore, the evaluation indicates that existing logs in various IoT-enhanced log formats can easily be translated into logs in our CORE metamodel format, demonstrating the achievement of the interoperability goal of the model.

### v2

@@ 6.3 Data Transformation Layer
The Data Transformation Layer focuses on transforming the CORE data structures into the OCEL 2.0 standard. This transformation step is critical, as it operationalizes the conceptual design (5) by aligning the internal semantics of the CORE objects, events, and relationships with the structural and syntactic constraints of OCEL 2.0. At its core, this layer transforms the CORE metamodel classes (Figure˜6) into OCEL classes (Figure˜2). This transformation is performed in three phases: The first phase transforms physical items, digital artifacts, and data sources from the CORE metamodel into OCEL-compatible objects. As illustrated in Figure˜12, the CORE implementation provides multiple utility classes. As described in Section˜5.2.2, Core Objects are grouped into three categories (i.e., DataSources, BusinessObject, GeneralObject). This subdivision can be refined using the object_class attribute of the object class. Moreover, the enum contains distinct possible object classes that are semantically connected to categorize. For example, a sensor is likely to be a data source. In addition to the object_class attributes, the object_type and object_id attributes are assigned directly to the corresponding OCEL counterparts. Additionally, all use case-specific attributes are assigned in the columns ‘ocel:attr:[key]’, where the key represents the corresponding attribute label. Finally, each object_id must be unique; otherwise, the object will not be transferred.
@@
The second phase handles the event transformation. For that, the implementation provides an abstract Event class with three implementations (see Section˜5.2.1, Figure˜13). The event class requires four attributes, while the ProcessEvent implementation requires one additional activity attribute. The event_id, timestamp, event_class and attributes are mapped accordingly to the object implementation. The event_type is dependent on the event_class. First, if the event_class is process_event the event_type is the activity label. Next, if the event_class is iot_event the event_type is some meaningful label. Finally, if the event_class is observation, the event_type is hardcoded to observed.
@@
The last phase establishes and maintains three relationships between the entities. As previously outlined, utility classes are also present in this phase (Figure˜14). They define the required attributes for the different relationships. Generally, the (event/object)_id is utilized for the mapping between the different entities. Each relationship is associated with a relationship_class, which is also serialized in the relationship table. As there is a direct equivalent in the OCEL 2.0 standard for the e2o and o2o relationships, these can be transferred directly. The only exception is the mandatory activity that must be used for each relationship (in OCEL). As this is not required for every event_class, the event_type is used in those cases. On the other hand, the event-to-event relationships require more parsing because there is not a corresponding definition in the OCEL standard. Given that, a linking object without semantically relevance is utilized. Therefore, two relationships are established. One between the first event and the linking object, and another between the linking object and the second event. Hence, the data can be queried as there are directly connected.
@@ 6.4 OCEL Representation Layer
The implementation should be independent of a specific OCEL implementation (e.g., pm4py.OCEL [(pm4py)], ocpa [(ocpa)]) and follow the structure provided in Figure˜2. This abstraction not only enhances the maintainability and flexibility of the implementation, but also ensures that the output remains reusable and interoperable.
@@ 7.1.1 Use case description
In this use case, we describe how we transcribe a log originally stored in the NICE format to our common CORE metamodel . This synthetic event log, first presented in (bertrand2023nice), represents the behavior of two inhabitants living in a smart home for four weeks. To generate it, we used the smart spaces simulator presented in (veneruso2023model), which we adapted to generate NICE logs. More specifically, the two users were programmed to execute daily life activities such as Cook and eat, Do the dishes, Drink, Eat cold, Eat warm, Exercise, Go work, Sleep, Rest, Shop, Use Computer, Watch tv and Wc. Moreover, the simulator was programmed so that during one week, the users show a different behavior, i.e., they stay at home the whole day instead of going to work in the morning and coming back home in the evening during weekdays. This change in behavior is due to extremely high outside temperatures during that week, which are tracked by temperature sensors and logged as IoT events.
@@
  • Flexibility of the process: human routines are characterized by their context dependency, their lack of structure, the repetition of activities and the execution of multiple activities in parallel (Leotta2015ApplyingPM);
@@ 7.2.1 Use Case Description
In manufacturing processes, multiple (IoT) devices operate collaboratively to produce a part. These processes encompass various steps, including the handling of parts, which may involve multiple machines, transportation between different machines, and the measuring of parts. In the following, two use cases are presented that demonstrate the integration of IoT devices into manufacturing processes. In each case, the processes are managed by a process execution engine and are logged using the XES DataStream extension, ensuring real-time tracking and analysis of production events. This approach facilitates enhanced process monitoring, optimization, and traceability, contributing to improved efficiency and product quality. (datastream):
  • In the IoT lab at Trier University [^5] More information can be found at https://iot.uni-trier.de., a Fischertechnik learning factory is used to carry out research related to BPM (Malburg.2020_BPMAndIoT). The use case includes machines which are equipped with several types of sensors that collect production-related data, such as the state of connected light barriers or switches. Complementing this Internet of Things (IoT) data, information pertaining to process executions is systematically recorded and maintained as event logs, facilitating comprehensive monitoring and analysis of manufacturing processes (Grueger.2023).
@@ 7.2.2 Translation to core log format
The logs for the use cases described before are logged using the XES DataStream extension (datastream). In the following, it is described how they can be translated into the core model format:
@@ 7.3.1 Use case description
In this use case, the focus is on the IoT-supported online monitoring and conformance checking of the blood donation process, augmented with hand hygiene guidelines from the World Health Organization (WHO) that prescribe additional process steps of hand disinfection (franceschetti2023proambition). The goal is to be able to provide a timely feedback to the healthcare workers regarding the conformance with the hand hygiene requirements, reducing the risk of contamination that could threaten the patients’ health. As the process is fully manual and executed without the support of a process engine, the challenge here is to be able to generate in real-time IoT events from the process execution and abstract them into the respective process events. However, ambiguities deriving from multiple process events such as Apply tourniquet and Insert needle being represented by the same IoT events pose a challenge to the event abstraction. This results in multiple process events being generated from a given set of IoT events, with attached a quantification of the respective ambiguity.
@@ 8 Discussion
The set of requirements presented in Sect. 4 is derived from the harmonization of requirements from different models addressing diverse use cases. As such, it provides a solid foundation for the development of a suitable representation of IoT-enhanced process logs. As reported in Sect. 7, the proposed CORE metamodel successfully fulfills these requirements, effectively supporting different use cases in various IoT domains. Moreover, the application of the proof-of-concept implementation to concrete use cases underscores the suitability of the metamodel for multiple near real-world applications.
The CORE metamodel offers the following implications for researchers:
  • Unified Representation: The model provides a standardized approach to integrating IoT data, process data, and contextual information across different levels of granularity (cf. Sect. 7). Thus, the CORE metamodel embodies a common reference point for future research, facilitating cross-study comparisons, meta-analyses, and the development of new process mining algorithms.
  • Flexibility and Extensibility: By supporting semantic annotations, metadata representation, and diverse data sources, the model accommodates a wide range of use cases that require the representation of additional data beyond events, as demonstrated in the evaluation in Sect. 7. This extensible architecture paves the way for future extensions and customized adaptations, for example capturing domain-specific knowledge. We expect this to be instrumental to enable further advancements in context-aware as well as semantics-aware process mining (rebmann2022enabling).
  • Interoperability: We deliberately made the choice of implementing the CORE metamodel within the OCEL 2.0 format (berti2024ocel). Currently, OCEL 2.0 is the go-to format adopted by researchers in process mining (goossens2024object). Therefore, we expect that this will facilitate uptake by researchers and integration with existing use cases, fostering data exchange and reuse.
@@
Moreover, the CORE metamodel yields the following implications for industry practitioners:
  • Data Exchange and Integration: The adoption of the CORE metamodel allows organizations to translate and integrate data from various sensors, information systems, and legacy event log formats. The definition of mappings between their existing IoT and process-related data and the CORE metamodel facilitates data exchange and integration between different systems.
  • Enhanced Process Analytics: The metamodel design allows for the simultaneous representation of process events and IoT observations. This duality enables practitioners to bridge the gap between the physical world and digital process representations. This enhances the ability to monitor and analyze business processes at different levels of granularity, thereby obtaining more sophisticated insights into enterprise operations.
@@
Overall, the CORE metamodel addresses several challenges, providing benefits for researchers and practitioners. Still, we recognize some limitations of the metamodel. Specifically, the reconciliation of diverse requirements from different use cases and research groups resulted in a number of trade-offs in the expressive power of the metamodel. This results in the lack of explicit concepts in the CORE metamodel able to represent use case-specific concepts, such as the Formal Model concept from CAIRO, which represents a formal process specification and is relevant in the smart healthcare use case. Nevertheless, the CORE metamodel remains open for specialization and extension to ensure that also use-case specific requirements can be met. This can be achieved, for example, by leveraging event attributes or by specializing the general object concept and its associations (cf. Sect. 5). This extensibility can be supported by a model-driven engineering (MDE) (kent2002model) approach, where the CORE metamodel acts as a central specification from which new artifacts can be derived. With this approach, adopting the metamodel and introducing new subclasses or domain-specific concepts does not require a full re-implementation, as changes can be integrated (semi-)automatically using MDE tools. We emphasize that these trade-offs reflect deliberate design choices made during the metamodel development to prioritize generalizability and interoperability–which we consider major goals of the metamodel.
@@ 9 Conclusion and Future Work
In conclusion, the contribution of this paper is manifold: 1) we present a set of requirements, important for the development of a suitable representation format for IoT-enhanced event logs, 2) we propose a new model referred to as CORE metamodel for IoT-enhanced event log, rooted in the object-centric paradigm to foster data exchange in the community, 3) we present our prototypical Python implementation of this new model, compatible with the PM4PY package and the OCEL 2.0 specification, and 4) we use this implementation for evaluation in several near real-world IoT use cases. To ensure the quality of the model and rigor in its design, the model has been created following the design science research methodology, building on existing IoT ontologies and IoT-enhanced log models, and based on a set of requirements gathered from experts in IoT-enhanced PM. Furthermore, the application of the metamodel to multiple use cases indicates that existing logs in various IoT-enhanced log formats can easily be translated into logs in our CORE metamodel format, demonstrating the achievement of the interoperability goal of the model.